# For Better or For Worse? Learning Minimum Variance Features With Label Augmentation

### v2

@@ 4.1 Spurious Pixels in Binary Classification
Figure 1: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.1.
@@
We first consider the case of logistic regression trained on data that has trivial spurious correlations that do not appear in the test data. For our binary classification task, we consider reductions of the standard CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) benchmarks to binary classification by fixing two classes from each dataset as the positive and negative classes, and replacing their original labels with the labels $+1$ and $-1$ respectively. This leads to 10000 training examples and 2000 test examples in the case of CIFAR-10, and 1000 training examples and 200 test examples in the case of CIFAR-100. Our experiments are not sensitive to the choice of binary reduction, and for the experiments in this section we fix the $-1$ class to be class 00 from the original data and the $+1$ class to be class $1$ from the original data (“airplane” and “automobile” for CIFAR-10, “apple” and “aquarium fish” for CIFAR-100).
@@
We preprocess the training data to have mean zero and variance one along every image channel (and apply the same transform to the test data). To introduce spurious correlations into the training data, we modify the data such that the first value in the tensor representation of each training input (corresponding to the “R” channel of the top left pixel in the original image) is replaced by $y\gamma$ with $\gamma=0.1$ (the choice of $\gamma$ here is not important as long as it is small, we verified the results for $\gamma=10^{-5}$ up to $\gamma=0.1$). This ensures that the training data is trivially linearly separable in the first dimension of the data, but learning this first dimension requires having larger weight norm due to the scaling by $\gamma$. We leave the test data unchanged - our goal is to determine whether models trained on the modified training data can learn more than just the single identifying dimension.
@@
We then train logistic regression models to convergence (100 epochs) on both the reduced CIFAR-10 and CIFAR-100 tasks across a range of settings for weight decay, label smoothing, and Mixup. We consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter is obtained from the experiments of Müller et al. (2020). For Mixup, we fix the mixing distribution to be the canonical choice of $\mathrm{Beta}(\alpha,\alpha)$ introduced by Zhang et al. (2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice. We train all models using Adam with the hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $5*10^{-3}$, and a batch size of 500.
After training, we report the mean test error along with one standard deviation error bounds over five training runs for each model setting in Figure 1. As can be seen from the results, both the label smoothing and Mixup models have high test error for all settings while weight decay achieves a significantly lower test error for all $\lambda>0$. The higher variance in the reported CIFAR-100 test errors can almost certainly be attributed to the fact that our binary version of CIFAR-100 consists of only 1000 training points and 200 points.
@@
4.2 Background Shift in Multi-Class Classification
(a) Shaded MNIST
@@ 4.2 Background Shift in Multi-Class Classification
Figure 3: Final model test errors over our hyperparameter sweep for the shaded MNIST dataset of Section 4.2, alongside a visualization of samples from the dataset.
Although the experiments of Section 4.1 cleanly verify our theory, the spurious correlation introduced in the training data is effectively adversarial in nature. We now consider the case of a spurious correlation of greater practical interest: backgrounds that are correlated with the correct class label.
In particular, we consider a very similar setup to the one used by Arjovsky et al. (2020) to motivate the influential invariant risk minimization framework: namely, we construct a shaded version of the MNIST dataset in which the background pixels for each class are replaced with different low intensity colors corresponding to the class labels. Unlike Arjovsky et al. (2020), however, we maintain all 10 classes since our theory from Section 3.2 suggests that our observations should generalize to the multi-class setting. The colors used for each class background are permuted between the train and the test data, so that a model that learns to predict only using background pixels cannot achieve good test accuracy.
The only constraints we place on the background colors are that they are distinct across classes and that their intensities (i.e. values in RGB space) are small (it suffices to consider values bounded by 16); this makes it difficult to visually distinguish the background colors (hence "shaded MNIST") even though they are obviously different in RGB space. The color intensity constraint is not necessary for the failure of label smoothing and Mixup, but is necessary for weight decay to succeed since it ensures that the higher variance feature (the actual digit in the foreground) is generally more separable (due to larger pixel intensities). Note that although there is no variance in the background color conditional on a class, there is significant variance in the actual pixels per class as the locations of the background pixels change across data points.
For our model setup, we also follow Arjovsky et al. (2020) and consider a simple 2-layer feedforward neural network with ReLU activations and a hidden layer size of $2048$, as this is sufficient to achieve good performance on MNIST and is a tiny enough model that we can efficiently do the same hyperparameter sweep from Section 4.1. We follow the same training procedure as Section 4.1, training until the training loss roughly converges (only 20 epochs are necessary).
@@ 5 Conclusion
In this work, we have shown that label augmentation strategies such as label smoothing and Mixup exhibit a variance minimization effect that can lead to learning low variance, spurious correlations in the data even when there exists a better but higher variance feature. The main practical takeaway from this is that label smoothing and Mixup do not necessarily always improve the robustness of models, and their failure to do so may be indicative of certain identifying features in the training data (i.e. adversarially introduced patches or background pixels).
Limitations. The main theoretical limitation of work is that our results for general models are only loss lower bounds and do not characterize optimization dynamics. Our experiments that such dynamics for feedforward networks would still reflect our observations, however, and it would be an interesting follow-up to study these dynamics using the techniques of Zou et al. (2023) and Chidambaram et al. (2023). On the empirical side, our main limitation is that we have demonstrated our theory using only small-scale experiments; a comprehensive empirical evaluation of label smoothing and Mixup in larger scale spurious correlations settings guided by our theory would be a natural avenue for future work.
Broader Impacts. Although label augmentation are used to train and finetune large-scale models, our results concerning them in this work have mostly been theoretical. As a result, we do not anticipate any direct misuse of our results or any broader harmful impacts.
Acknowledgements
Rong Ge and Muthu Chidambaram are supported by NSF Award DMS-2031849 and CCF-1845171 (CAREER). Muthu would like to thank Annie Tang for thoughtful feedback during the early stages of this project.
@@ A.1 Helper Lemma
Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a twice-differentiable convex function satisfying $\gamma_{1}\mathrm{Id}\prec\nabla^{2}\phi\prec\gamma_{2}\mathrm{Id}$. Then for any square integrable random variable $X$ on $\mathbb{R}^{d}$ it follows that:
@@
Let us decompose the unique minimizer $w^{*}$ as $w^{*}=\alpha_{1}u^{*}+\alpha_{2}v^{*}$, where $u^{*}$ and $v^{*}$ are orthonormal and satisfy $u^{*}_{\mathcal{H}}=0,v^{*}_{\mathcal{L}}=0$ (i.e. $u^{*}$ is a normalized version of the $\mathcal{L}$ components of $w^{*}$ and $v^{*}$ is the same but for the $\mathcal{H}$ components). We claim that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e. $(x,y)$. Indeed, if this were not the case then by the assumption on $\pi$ we could choose an orthonormal vector $z^{*}$ with $z^{*}_{\mathcal{L}}=0$ that satisfies $y\left\langle z^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e. $(x,y)$ and decrease the loss of $w^{*}$ by replacing $v^{*}$ with $z^{*}$.
Now suppose that $\alpha_{1}>\alpha_{2}$. Then we claim that replacing $\alpha_{1}$ and $\alpha_{2}$ with $\gamma=\sqrt{(\alpha_{1}^{2}+\alpha_{2}^{2})/2}$ yields a solution with lower loss than $w^{*}$.
To see this, first observe that $2\gamma^{2}=\alpha_{1}^{2}+\alpha_{2}^{2}$, so that the norm of the modified solution is the same as $w^{*}$. This implies that the weight decay penalty term in $\ell_{\beta}$ is unchanged.
Furthermore, we have that $\gamma\in(\alpha_{2},\alpha_{1})$, and that $\gamma-\alpha_{2}>\alpha_{1}-\gamma$. This follows from the fact that $\gamma\geq(\alpha_{1}+\alpha_{2})/2$ by construction. This, combined with the fact that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$, then implies for $\pi$-a.e. $(x,y)$ that:
$y\left\langle\gamma u^{*}+\gamma v^{*},x\right\rangle>y\left\langle\alpha_{1}u^{*}+\alpha_{2}v^{*},x\right\rangle.$ (A.2)
Which contradicts the optimality of $w^{*}$. Therefore, we must have $\alpha_{1}\leq\alpha_{2}$, which gives the desired result. ∎
@@
The result follows from the same decomposition approach of Theorem 3.4 and the definition of the max-margin solution (Definition 2.1). Namely, we can move weight from the max-margin direction in $\mathcal{L}$ to the max-margin direction in $\mathcal{H}$ to decrease the norm while satisfying the separation constraint. ∎
@@
Once again, since we can restrict our attention to $g(X)\in[\gamma,1-\gamma]$ for some $\gamma$ (as the loss goes to infinity if $g(X)=1$ on a set of positive $\pi_{X}$-measure with $\alpha>0$), it is easy to verify that (A.12) is strongly convex in $g(X)$. The desired result then follows by applying Lemma A.1 with the regular conditional distribution $g(X)\mid Y=y$ for each term in (A.13). ∎

### v3

@@ 4.2.1 Binary Classification With Perturbed Training Data
Figure 2: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.2.1.
@@
For our binary classification task, we consider reductions of the CIFAR benchmarks to binary classification by fixing two classes from each dataset as the positive and negative classes, and replacing their original labels with the labels $+1$ and $-1$ respectively. Our experiments are not sensitive to the choice of binary reduction, and for the experiments in this section we fix the $-1$ class to be class 00 from the original data and the $+1$ class to be class $1$ from the original data.
@@
We preprocess the training data to have mean zero and variance one along every image channel, but do not use other augmentations or perform feature extraction using pretrained models. We then “adversarially” modify the training data such that the first value in the tensor representation of each training input is replaced by $\gamma y$ with $\gamma=0.1$ ($\gamma$ here just needs to be small, we verified the results for $\gamma=10^{-5}$ up to $\gamma=0.1$). This ensures that the training data is linearly separable in the first dimension of the data, but learning this first dimension requires having larger weight norm due to the scaling by $\gamma$. We leave the test data unchanged - our goal is to determine whether models trained on the modified training data can learn more than just the single identifying dimension.
@@
We then train logistic regression models on both the reduced CIFAR-10 and CIFAR-100 tasks across a range of settings for weight decay, label smoothing, and Mixup. We consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter is obtained from the experiments of Müller et al. (2020). For Mixup, we fix the mixing distribution to be the canonical choice of $\mathrm{Beta}(\alpha,\alpha)$ introduced by Zhang et al. (2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice. Other hyperparameters are the same as before, except we use a learning rate of $5\times 10^{-3}$.
The results across hyperparameter values are shown in Figure 2. Both the label smoothing and Mixup models have high test error for all settings while weight decay achieves a significantly lower test error for all $\lambda>0$. ERM also has high test error; in this case we differ from the setting of implicit bias results due to training with Adam and likely not training for the period required for convergence to the max-margin solution. Furthermore, these results are insensitive to introducing a small amount of weight decay to the label smoothing and Mixup models (i.e. $5\times 10^{-4}$ as in the previous subsection).
@@
4.3 Spurious Background Correlations
(a) Colored MNIST
@@ 4.3 Spurious Background Correlations
Figure 4: Final model test errors over our hyperparameter sweep for the colored MNIST dataset of Section 4.3, alongside a visualization of samples from the dataset.
For our multi-class analogue to Section 4.2.1, we consider a similar setup to the one used by Arjovsky et al. (2020) to motivate the influential invariant risk minimization framework: namely, we construct a colored version of the MNIST dataset in which the background pixels for each class are replaced with different colors corresponding to the class labels. Unlike Arjovsky et al. (2020), however, we maintain all 10 classes since our theory from Section 3.2 suggests that our observations should generalize to the multi-class setting. The colors used for each class background are permuted between the train and the test data, so that a model that learns to predict only using background pixels cannot achieve good test accuracy.
The only constraints we place on the background colors are that they are distinct across classes and that their intensities (i.e. values in RGB space) are small (it suffices to consider values bounded by 16). The latter constraint is not necessary for the failure of label smoothing and Mixup, but is necessary for weight decay to succeed since it ensures that the higher variance feature (the actual digit in the foreground) is generally more separable (due to larger pixel intensities). Note that although there is no variance in the background color conditional on a class, there is significant variance in the actual pixels per class as the locations of the background pixels change across data points.
For our model setup, we also follow Arjovsky et al. (2020) and consider a simple 2-layer feedforward neural network with ReLU activations and a hidden layer size of $2048$, as this is sufficient to achieve good performance on MNIST and is a tiny enough model that we can efficiently do the same hyperparameter sweep from Section 4.2.1. Training details remain the same as in Section 4.2.1, except we only train for 20 epochs as this is sufficient for the training loss to roughly converge and greatly expedites the hyperparameter sweep.
@@ 5 Conclusion
In this work, we have shown that label augmentation strategies such as label smoothing and Mixup exhibit a variance minimization effect (both in theory and in practice) that leads to lower variance intermediate representations and model outputs, which then correlate with better test performance on image classification benchmarks. A natural follow-up direction to our results is to investigate whether regularizers for encouraging lower variance features can be implemented directly to improve model performance, which would shed light on whether there is a causal relationship between this phenomenon and improved performance.
Ethics Statement
Although label smoothing and Mixup are used to train and fine-tune large-scale models, our results concerning them in this work have mostly been theoretical and explanatory. As a result, we do not anticipate any direct misuse of our results or any broader harmful impacts.
Reproducibility Statement
All proofs of the results in this paper can be found in Appendix A. The supplementary material contains the code necessary to generate all figures, with instructions on how to run each experiment.
Acknowledgments
Rong Ge and Muthu Chidambaram were supported by NSF Award DMS-2031849 and CCF-1845171 (CAREER) during the completion of this work. Muthu would like to thank Annie Tang for thoughtful feedback during the early stages of this project.
@@ A.1 Helper Lemma
Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a twice-differentiable convex function satisfying $\gamma_{1}\mathrm{I}_{d}\prec\nabla^{2}\phi\prec\gamma_{2}\mathrm{I}_{d}$. Then for any square integrable random variable $X$ on $\mathbb{R}^{d}$ it follows that:
@@
Let us decompose the unique minimizer $w^{*}$ as $w^{*}=\alpha u^{*}+\beta v^{*}$, where $u^{*}$ and $v^{*}$ are orthonormal and satisfy $u^{*}_{\mathcal{H}}=0,v^{*}_{\mathcal{L}}=0$ (i.e. $u^{*}$ is a normalized version of the $\mathcal{L}$ components of $w^{*}$ and $v^{*}$ is the same but for the $\mathcal{H}$ components). We claim that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e. $(x,y)$. Indeed, if this were not the case then by the assumption on $\pi$ we could choose an orthonormal vector $z^{*}$ with $z^{*}_{\mathcal{L}}=0$ that satisfies $y\left\langle z^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e. $(x,y)$ and decrease the loss of $w^{*}$ by replacing $v^{*}$ with $z^{*}$.
Now suppose that $\alpha>\beta$. Then we claim that replacing $\alpha$ and $\beta$ with $\gamma=\sqrt{(\alpha^{2}+\beta^{2})/2}$ yields a solution with lower loss than $w^{*}$.
To see this, first observe that $2\gamma^{2}=\alpha^{2}+\beta^{2}$, so that the norm of the modified solution is the same as $w^{*}$. This implies that the weight decay penalty term in $\ell_{\beta}$ is unchanged.
Furthermore, we have that $\gamma\in(\beta,\alpha)$, and that $\gamma-\beta>\alpha-\gamma$. This follows from the fact that $\gamma\geq(\alpha+\beta)/2$ by construction. This, combined with the fact that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$, then implies for $\pi$-a.e. $(x,y)$ that:
$y\left\langle\gamma u^{*}+\gamma v^{*},x\right\rangle>y\left\langle\alpha u^{*}+\beta v^{*},x\right\rangle.$ (A.2)
Which contradicts the minimality of $w^{*}$. Therefore, we must have $\alpha\leq\beta$, which gives the desired result. ∎
@@
We can apply the same decomposition as in the proof of Theorem 3.4; namely, $w^{*}=\alpha u^{*}+\beta v^{*}$ where $w^{*}$ is the max-margin solution and $u^{*}$ and $v^{*}$ are as before. Suppose again that $\alpha>\beta$ and let $\gamma=\sqrt{(\alpha^{2}+\beta^{2})/2}$. We claim that there exists $\epsilon\in(0,\gamma)$ such that $w=(\gamma-\epsilon)u^{*}+\gamma v^{*}$ satisfies $y\left\langle w,x,\geq\right\rangle 1$ for $\pi$-a.e. $(x,y)$, which would contradict $w^{*}$ being the max-margin solution since $\norm{w}^{2}<\norm{w^{*}}^{2}$.
From the exact same logic as in the proof of Theorem 3.4, we have:
$y\left\langle\gamma u^{*}+\gamma v^{*},x\right\rangle-y\left\langle w^{*},x\right\rangle>0$ (A.3)
for $\pi$-a.e. $(x,y)$. Since $\mathrm{supp}(\pi)$ is compact, (A.3) attains a minimum $\kappa>0$ over $\mathrm{supp}(\pi)$. Similarly, $y\left\langle u^{*},x\right\rangle$ attains a finite maximum. By choosing $\epsilon$ such that $y\left\langle\epsilon u^{*},x\right\rangle<\kappa$, we obtain the desired contradiction. ∎
@@
Once again, since we can restrict our attention to $g(X)\in[\gamma,1-\gamma]$ for some $\gamma$ (as the loss goes to infinity if $g(X)=1$ on a set of positive $\pi_{X}$-measure with $\alpha>0$), it is easy to verify that (A.13) is strongly convex in $g(X)$. The desired result then follows by applying Lemma A.1 with the regular conditional distribution $\pi_{X\mid Y=y}$ for each term in (A.14). ∎